### v2

# Augmenting Math Word Problems via Iterative Question Composing

Abstract
Despite recent progress in improving the mathematical reasoning ability of large language models(LLMs), solving competition-level math problems without the use of external tools remains challenging for open-source LLMs. In this work, we introduce the MMIQC dataset, a mixture of processed web data and synthetic question-response pairs, to equip base models with better mathematical reasoning skills. In different model sizes, the models fine-tuned on MMIQC consistently outperform their counterparts by a clear margin on MATH test set. Notably, DeepSeek-67B-MMIQC achieves a 41.0% accuracy, 4.2% higher than the previous open-source SOTA. Our experiments also show that a large part of the improvement can be attributed to our novel augmentation method IQC(Iterative Question Composing), where we iteratively ask an LLM to compose new questions from the given seed problems and do rejection sampling from another LLM. We have released MMIQC on this HuggingFace hub page.
Machine Learning, Large Language Models
1 Introduction
Although large language models have been demonstrated to be powerful in various applications(Chen et al., 2021; Brown et al., 2020; Ouyang et al., 2022; Park et al., 2023; Huang et al., 2022b), solving math problems that require complex reasoning skills remains a challenging task. On MATH(Hendrycks et al., 2021b), a competition level math problem benchmark, open-source base LLMs such as the Llama family(Touvron et al., 2023a, b) fail to correctly answer most of the problems.
Previous work tries to enhance the mathematical reasoning abilities of base models through fine-tuning them on domain-specific data. Specifically, One line of work (Azerbayev et al., 2023; Lewkowycz et al., 2022) collect math corpora from web and fine-tune the models on them, which is also known as the procedure of continued pre-training. Another line of work focuses on constructing synthetic data through rejection sampling (Yuan et al., 2023), distilling from GPT-4/GPT-3.5(Yue et al., 2023) or question bootstrapping(Yu et al., 2023), and then use the generated question-response pairs to perform supervised fine-tuning in the way described in (Taori et al., 2023; Ouyang et al., 2022). However, there still exists a large performance gap between these fine-tuned models and the most advanced close-source models such as GPT-4(OpenAI, 2023) and Gemini-Ultra(Team et al., 2023). Given that simply adding more data does not always lead to better performance as shown in (Yu et al., 2023), how to bridge the gap remains an open challenge.
This work tackles the challenge by combining the two lines of work. On one hand, we reuse the high quality corpora used in the pre-training stage during fine-tuning. Specifically, MMIQC contains around 1200k question-response pairs we filter and pre-process from the math.stackexchange.com pages included in (Computer, 2023). On the other hand, for the synthetic data part of MMIQC, we increase the diversity by using multiple kinds of augmentation methods listed below:
  • Prompting GPT-4 with an integrated version of the question bootstrapping prompts used in (Yu et al., 2023), and do rejection sampling with GPT-3.5-Turbo on both seed and augmented problems.
  • Using a modified prompt present in (Liu et al., 2023) to ask GPT-4 to generate similar problems with answers given seed problems of the training set of MATH. Although the generated answers can be wrong, we perform rejection sampling on these problems as well.
  • Performing IQC(Iterative Question Composing) with 4 iterations in total. We iteratively ask GPT-4 to compose new questions from the given seed problems and do rejection sampling with GPT-3.5-turbo.
  • Filtering a 204k subset of MetaMathQA(Yu et al., 2023) and adding it to MMIQC.
More details of MMIQC will be introduced in Section 3.2.
Figure 1: The performance of base models and their fine-tuned versions on MATH benchmark. The results corresponding to the models remarked with an $*{}^{*}$ are evaluated by us. We can see that the models fine-tuned on MMIQC consistently outperform their counterparts by a clear margin.
We fine-tune some base models on MMIQC, resulting in models consistently achieve a large margin compared to their counterparts when evaluated on MATH, as shown in Figure 1. Specifically, Mistral-7B-MMIQC obtained by fine-tuning Mistral-7B(Jiang et al., 2023) on MMIQC, Llemma-34B-MMIQC for Llemma-34B(Azerbayev et al., 2023) and DeepSeek-67B-MMIQC for (Bi et al., 2024) achieve 36.0%, 38.6% and 41.0% accuracy on MATH, 5.8%, 3.8% and 4.2% higher than the previous SOTAs in their model size levels, respectively.
We highlight our contributions as follows:
  • We propose IQC(Iterative Question Composing), an augmentation method that can iteratively generate diverse data starting from a seed dataset of math word problems.
  • We release MMIQC, a mixture of processed web data and synthetic question-response pairs. In different model sizes, the models fine-tuned on MMIQC consistently outperform their counterparts by a clear margin on MATH test set. Notably, DeepSeek-67B-MMIQC achieves a 41.0% accuracy, 4.2% higher than the previous open-source SOTA.
  • Our results show that reusing the high-quality data in the pre-training corpora during the fine-tuning stage can improve the model performance, successfully combining the two lines of work of continued pre-training and supervised fine-tuning.
  • Our results also show that using multiple augmentation methods to construct datasets for fine-tuning is an efficient way to boost the performance of LLMs.
2 Related Work
Fine-tuning Base LLMs on Mathematical Datasets. Current practice of fine-tuning base LLMs on mathematical datasets can be classified into two kinds: 1) continued pretraining(Lewkowycz et al., 2022; Azerbayev et al., 2023). This line of work typically collects billion-tokens level mathematical text data from web and fine-tune the model in a way similar to that in the pre-training stage. 2) SFT(Supervised Fine-Tuning)(Yuan et al., 2023; Yu et al., 2023; Yue et al., 2023; Gou et al., 2023). Works in this line collect question-response pairs via various methods and train the models on their dataset in an Alpaca style. Our work locates in the mid between these two: MMIQC is a mixture of pre-training corpus and question-response pairs.
Reasoning Frameworks for Solving Mathematical Problems. Much effort has been devoted to achieving a higher accuracy on math word problems benchmarks by designing different procedures of using the given LLMs to obtain the answers, which we refer to as reasoning frameworks. Among them, Prompting-based methods (Radford et al., 2019; Wei et al., 2023; Fu et al., 2022) play a significant role in activating the potential reasoning abilities for base LLMs through carefully designing the prompts shown to the models. Self-consistency(Wang et al., 2023b) samples multiple rationale paths for a model and then selects the answer by majority voting. Moreover, (Lightman et al., 2023; Wang et al., 2023a) use verifiers to help select the answer with highest score from the sampled answers of LLMs. Some frameworks also include the use of plug-in tools and external APIs. Program-aided prompting(Gao et al., 2022; Yue et al., 2023) provides in-context samples containing python codes for LLMs and use code interpreters to execute the output to facilitate reasoning. (Gou et al., 2023) interleaves natural language rationales with Sympy code and finetunes the model to follow their framework. We note that our results in Figure 1 do not include multiple times of sampling, use of verifiers or code interpreters, thus cannot be directly compared with the results reported in these works.
3 Method
In this section, we first state our proposed IQC (Iterative Question Composing) method in Section 3.1, and then introduce the make-up of MMIQC in Section 3.2 in detail.
3.1 Iterative Question Composing
Algorithm 1 Iterative Question Composing
1:  Require: question composing model $\pi_{q}$, rejection sampling model $\pi_{2}$, answer extractor that defines $\simeq$, text templater $x(\cdot,\cdot)$ with an inverse $x^{-1}(\cdot)$, initial seed dataset $S_{0}=\{(q_{i},a_{i})\}_{i=1}^{n}$, total #iterations $K$, question composing prompts $p_{1},p_{2},\cdots,p_{K}$, rejection sampling prompt $p_{r}$, maximal number of rejection sampling for one problem $m$
2:  for $k=1$ to $K$ do
3:     $S_{k}\leftarrow\{\}$
4:     $R_{k}\leftarrow\{\}$
5:     for all $(q,a)\in S_{k-1}$ do
6:        $x^{\prime}\sim\pi_{q}\left(\cdot|p_{k}\oplus x(q,a)\right)$
7:        $(q^{\prime},a^{\prime})\leftarrow x^{-1}(x^{\prime})$
8:        $S_{k}\leftarrow S_{k}\cup\{(q^{\prime},a^{\prime})\}$
9:        for $j=1$ to $m$ do
10:           $a^{(j)}\sim\pi_{r}(\cdot|p_{r}\oplus q^{\prime})$
11:           if $a^{(j)}\simeq a^{\prime}$  then
12:              $R_{k}\leftarrow R_{k}\cup\{q^{\prime},a^{(j)}\}$
13:           end if
14:        end for
15:     end for
16:     $D_{k}\leftarrow S_{k}\cup R_{k}$
17:  end for
18:  Output: $D_{1},D_{2},\cdots,D_{K}$
We first introduce some notations. To keep our statement brief, we refer to a combination of an LLM, its corresponding tokenizer and encoding/decoding method, and a fixed generation configuration (including the type of generation strategy, the sampling temperature and the stopping words, etc.) simply as ‘an LLM’ below. For an LLM $\pi$, we use $\pi(\cdot|p)$ to denote the distribution of the output of the LLM given the prompt $p\in\mathcal{A}^{*}$. For two paragraph of texts $p_{1}$ and $p_{2}$, we use $p_{1}\oplus p_{2}$ to denote the concatenation of them.
To perform IQC, we shall first specify an LLM $\pi_{q}$ for question composing, and another model $\pi_{r}$ for rejection sampling. For rejection sampling, we need an answer extractor to extract answer from a response. For two responses $r_{1},r_{2}$, we say $r_{1}\simeq r_{2}$ if we can extract a same answer from them using the pre-defined answer extractor. We also need a seed dataset $S_{0}=\{(q_{i},a_{i})\}_{i=1}^{n}$ of $n$ question-response samples (e.g. the training set of MATH) to start with.
In iteration #1, we prompt $\pi_{1}$ with $p_{1}\oplus x(q,a)$ for each $(q,a)\in S_{0}$, where $x(\cdot,\cdot)$ is a text templater that transforms a question-response pair into plain text with an inverse $x^{-1}(\cdot)$, and $p_{1}$ is a prompt that asks the model to compose a new question (with answer) in the same format with $x=x(q,a)$. Thus, we get a new dataset
$S_{1}=\{(q^{\prime}_{i},a^{\prime}_{i})\}_{i=1}^{n},$
with $(q^{\prime}_{i},a^{\prime}_{i})=x^{-1}(x^{\prime}_{i})$, where $x^{\prime}_{i}\sim\pi_{q}\left(\cdot|p_{1}\oplus x_{i}\right)$ is the output of question composing model for the $i$th sample. To maximally utilize the generated questions, we rejection-sample from $\pi_{r}$ on $S_{1}$ to get
$R_{1}:=\{(q^{\prime}_{i},a^{(j)}_{i})|a^{(j)}_{i}\simeq a^{\prime}_{i},i\in[n],j\in[m]\},$
with $a^{(j)}_{i}$ being the responses i.i.d. sampled from $\pi_{r}(\cdot|p_{r}\oplus q^{\prime}_{i})$, where $p_{r}$ is a prompt that asks the model to respond in a format fit for the answer extractor. The dataset obtained by uniting $S_{1}$ and $R_{1}$,
$D_{1}:=S_{1}\cup R_{1},$
is the desired IQC Iter #1 dataset.
In iteration #$k$, one simply repeats the procedure stated above with $S_{k-1}$ as the seed dataset. It is possible for one to take different question composing prompts $p_{k}$ in different iterations. The whole procedure of IQC is summarized in Algorithm 1.
Table 1: The make-up of MMIQC.
Data # Samples # Epochs Ratio
MetaMathQA 203.7K 3 26.6%
AnsAug & QB 66.5K 3 8.7%
AugSimilar 38.2K 3 5.0%
IQC 55.1K 3 7.2%
MathStEx 1203.6K 1 52.5%
3.2 MMIQC
You will be provided with a math problem in newline-delimited json format. Please augment 5 diverse problems from the given problem. The way you augment a problem can be: - Rephrase the problem. - Change the scenario without modifying specific quantities. - Set 1 number in the problem to an unknown variable, put the answer in the problem and ask what is the value of the variable. Ensure the generated problem is reasonable. Otherwise, skip this method. - Other approaches that can ensure the correctness of the answer you provide to the augmented problem. Your response should only contain text in newline-delimited json format, keeping the same with the given problem. Please use two backslashes to represent one in the strings.
Figure 2: The prompt we use to perform question bootstrapping for GPT-4.
You will be provided with 1 math problem in newline-delimited json format. Please generate 3 diverse new problems similar to the given problem. Your response should only contain text in newline-delimited json format, keeping the same with the given problem. The solutions to the generated problems should be as brief as possible. Ensure there is only one box in the solution and the answer is completely the same with the content in the box. Please use two backslashes to represent one in the strings.
Figure 3: The prompt we use to generate questions similar to the seed problems for GPT-4.
[…]
Figure 4: The prompt we use to perform question composing in IQC. The brown part is not used in iteration #1.
In this subsection, we introduce how each part of MMIQC is constructed in detail.
Subset of MetaMathQA. The original MetaMathQA dataset is constructed by sampling GPT-3.5 for $k=20$ times under a $T=0.7$ temperature for each problem in the training set of MATH/GSM8K, or its bootstrapped versions. We restrict the number of samples for each completely same question to be 3 and 1 for MATH and GSM8K, respectively, to obtain a subset of MetaMathQA. This subset contains 112.2K GSM8K question-response pairs and 91.5K MATH pairs.
Answer Augmentation and Question Bootstrapping. We integrate the question bootstrapping methods used in (Yu et al., 2023) into a single prompt shown in Figure 2. Our motivation is that, given GPT-4 is a highly natural language understanding, a few-shot prompting style used in (Yu et al., 2023) might suppress the diversity of the augmented questions. The seed dataset is the set of samples in the training set of MATH that do not contain Asymptote language in their question statements. We rejection sample from GPT-3.5 on both the seed dataset and generated questions, obtaining 66.5K question-response pairs. We use a temperature $T=1.0$ for both question bootstrapping and rejection sampling.
Augmented Similar Problems. (Liu et al., 2023) asks GPT-3.5 to generate 10 similar questions given 1 seed problem in the training set of GSM8K. In our practice, we find that GPT tends to generate several almost the same problems regardless of the given seed problem when asked to generate up to 10 new problems. Thus, we only ask GPT to generate 3 problem (with a solution, for rejection sampling) each time, using the prompt in Figure 3. Considering rejection sampling needs the answer to the problem better to be correct, we use the stronger GPT-4 instead of GPT-3.5. In order to control the cost, our prompt emphasizes that the solution should be as brief as possible. The total number of the augmented similar problems and the question-response pairs rejection sampled from them is 38.2K. We use a temperature $T=1.0$ for both procedures.
Iterative Question Composing. We perform Iterative Question Composing for 4 iterations as described in Section 3.1. Specifically, we use GPT-4 for question composing model $\pi_{q}$ with a $T=0.7$ temperature and GPT-3.5 for rejection sampling model $\pi_{r}$ with a $T=1.0$ temperature. The question composing prompts are shown in Figure 4. The text templater $x(\cdot,\cdot)$ we use is to transform a question-response pair into JSON text format, with fields ‘problem’ and ‘solution’. The resulting dataset has 55.1K samples in total. [^2] A part of the samples are generated by performing IQC for 2 iterations using a legacy version of prompts.
Mathematics Stack Exchange. We observe that in OpenWebMath(Paster et al., 2023) dataset, the data from Mathematics Stack Exchange shows high quality and is most related to competition level math. Motivated by this, we extract the data collected from Mathematics Stack Exchange in RedPajama(Computer, 2023) and pre-process it into question-response pairs. For each Mathematics Stack Exchange page, we only retain the answer with highest votes. Then we filter out the answer that does not contain a formula environment symbol ‘\$’. This results in a dataset with 1203.6K question-response pairs.
Table 1 shows the make-up of MMIQC. MMIQC contains 3 epochs of the subsets mentioned above, except for the Mathematics Stack Exchange part.
4 Experiments
Table 2: A comparison of the accuracy of our fine-tuned models and existing models on MATH. Our evaluation result is marked with an ***. Other results, unless specified with an citation, are from (Wang et al., 2023a).
Model FT-Dataset Tool Usage? Eval Method MATH(%)
Minerva-540B (Uesato et al., 2022) Arxiv+Web No maj1@64 50.3
GPT-4 Unknown No pass@1 42.4
GPT-4-0613 Unknown No pass@1 56.2
Gemini-Ultra(Team et al., 2023) Unknown No pass@1 53.2
Llama-2-7B(Touvron et al., 2023b) N/A No pass@1 2.5
Llemma-7B (Azerbayev et al., 2023) Proof-Pile-2 No pass@1 18.0
MetaMath-7B(Yu et al., 2023) MetaMathQA No pass@1 19.8
Mistral-7B-MetaMathQA(Yu et al., 2023) MetaMathQA No pass@1 28.2
Mistral-7B-MMIQC* MMIQC No pass@1 36.0
MAmmoTH-Coder-7B(Yue et al., 2023) MathInstruct Code pass@1 35.2
ToRA-Code-7B (Gou et al., 2023) ToRA-Corpus Code pass@1 44.6
CodeLlamma-34B N/A Code pass@1 25.0
Llemma-34B-MetaMathQA MetaMathQA No pass@1 34.8
Llemma-34B-MMIQC* MMIQC No pass@1 38.6
Llemma-34B-MetaMathQA MetaMathQA Shepherd maj+verify1@256 47.3
ToRA-Code-34B (Gou et al., 2023) ToRA-Corpus Code pass@1 50.8
DeepSeek-67B(Bi et al., 2024) N/A No Pass@1 18.7
Deepseek-67B-MetaMathQA MetaMathQA No pass@1 36.8
Deepseek-67B-MMIQC* MMIQC No pass@1 41.0
Deepseek-67B-MetaMathQA MetaMathQA No maj1@256 45.4
Deepseek-67B-MetaMathQA MetaMathQA Shepherd maj+verify1@256 48.1
4.1 Fine-tuning Setup
Our fine-tuning strategy mainly follows the practice of (Taori et al., 2023), except that we use a different prompt template to transform the question-response pairs. For a sample from Mathematics Stack Exchange, the corresponding prompt fed into the model during training is a simple concatenation of the question and response with two new-line symbols. For a sample from other subsets, we additionally add a prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’ to the question-response concatenation.
We use the HuggingFace transformers library (Wolf et al., 2019) for our fine-tuning experiments. We fine-tune all models on MMIQC for 1 epoch, using a $3\%$ warmup ratio linear learning rate schedule with a maximum learning rate $51e-$ 5. Employing the DeepSpeed Zero-3 Stage (Rajbhandari et al., 2020), we finetune 7B models on one node of 8xA800 gpus with micro batch-size at 8 and gradient accumulation at 4, 34B models on 2 nodes with micro batch-size at 4 and gradient accumulation at 4, maintaining an effective batch-size of 256.
4.2 Model Evaluation
We evaluate the fine-tuned models on MATH(Hendrycks et al., 2021a), a competition-level math word problems benchmark with 5000 test problems. We prompt all our fine-tuned models with the test question with the prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’, and extract the answer from the output using a modified version of the answer extractor provided in (Lewkowycz et al., 2022). We use a series of rules to infer whether the extracted answer is the same with the ground-truth answer, including a comparison using SymPy(Meurer et al., 2017).
The complete results of our evaluation and a comparison with existing models are shown in Table 2.
4.3 Ablation Study of Subsets of MMIQC
Table 3: How different subsets of MMIQC affect the accuracy of the finetuned model on MATH. The base model is Mistral-7B, and the accuracy baseline is the reported 28.2% of Mistral-7B-MetaMathQA.
Data # Samples MATH(%)
MetaMathQA(subset) 203.7K 26.4(-1.8)
+ AnsAug & QB 270.2K 30.1(+1.9)
+ AugSimilar 308.4K 31.5(+3.3)
+ IQC iter #1 330.2K 33.0(+4.8)
+ IQC iter #2 346.2K 33.7(+5.5)
+ IQC iter #3&#4 363.5K 34.4 (+6.2)
+ MathStackExchange +1203.6K 36.0 (+7.8)
In order to understand the ratio of contribution to the improvement revealed in Table 2 of different subsets of MMIQC, we fine-tune Mistral-7B with a series of training sets constructed by gradually adding the subsets. When MathStackExchange is not added, we fine-tune for 3 epochs. It can be seen from Table 3 that
  • Although our filtered subset of MetaMathQA is only half the size of the original dataset (which has 395K samples, more than the total number of samples of our synthetic data), the performance drop is only 1.8%. This shows that the $k=20$ strategy in (Yu et al., 2023) results in some redundancy.
  • Our Answer Augmentation & Question Boosting data help the fine-tuned model beat Mistral-7B-MetaMathQA, verifying our hypothesis that directly asking GPTs to perform question bootstrapping is more efficient than providing few-shot examples to them.
  • Our IQC method leads to a significant 3.1% improvement from a high accuracy $31.5\%$ with only 55.1K samples, showing its efficientcy. Moreover, the later iterations of IQC also account for a certain ratio of improvement, proving that IQC is a method that can continuously generate new data that can help increase the diversity when added to the data generated in previous iterations.
4.4 Contamination test
We check the $n$-gram matches for MMIQC to ensure that the improvement is not a result of direct memorization. We use the script provided by (Azerbayev et al., 2023) to check the $n$-gram matches between the synthetic part of the MMIQC and MATH test set. It turns out that for a 30-gram match check, there are 44 hits of match between the ‘solution’ field of MATH test set and the ‘output’ field of MMIQC, far fewer than the 168 hits between that of MATH test set and MATH training set. Moreover, we manually check these 44 hits and find that 43 among them belong to the case where intermediate steps of the solutions to similar but different questions collide, with the only exception being the question ‘A regular polygon has interior angles of 144 degrees. How many sides does the polygon have?’. This almost rules out the possibility that fine-tuned models get memorization of solutions to the problems in the test set, indicating a very low risk of data contamination for MMIQC.
5 Conclusion
In this work, we introduce a novel data augmentation method for math word problem datasets called IQC and use it in the construction of our MMIQC dataset. Our evaluation results show that the models fine-tuned on MMIQC achieve new SOTAs on the MATH benchmark. The improvements of our models benefit from the diverse data source of MMIQC and the effectiveness of IQC. For future directions, we are interested in how to equip open-source models with the ability of composing questions, in order to perform IQC in a self-evolution style, similar to that in (Huang et al., 2022a). Besides, how to integrate the verification systems of (Wang et al., 2023a; Liu et al., 2023), which are originally used to improve the accuracy during inference time, into the procedure of IQC, is also an attractive topic.